An object is launched with an initial velocity of 10 meters per second from the ground level at an angle of 53° above the horizontal. What is the horizontal range of the object? (cos53°=0,6 sin 53°=0,8).
Take g=10 m/s2
1
Expert's answer
2020-10-30T13:18:52-0400

"s=\frac{v^2sin2a}{g}=\frac{2v^2sinacosa}{g}"

"s=\frac{2\cdot10^2\cdot0.6\cdot0.8}{10}=9.6" (m).